AGw Commissioning with IRTC - 14-15 November 2008 MST

Observers: A. Rakich, J. Hill, N. Ageorges in Tucson
Telescope Operator: W. Wack

Summary

Ran acquisition tests to try to understand acquisition failures seen in recent nights.

Took some data for Jesper for measuring AGw coordinate transformations, but seeing was very poor.

We exercised lots of presets, offsets and dithers as LUCIFER, and as IRAF LBTtools acting as LUCIFER. No significant issues were found.

Took some field aberration data in seeing that had improved to merely bad.

Finally got some more data for Jesper on M67 in decent seeing at the end of the night.

Details

Open

6:12 usual startup problems. Right shutter door not opening, mucked around a bit finding a suitable BS star at elevation, preset to BS9188

6:32 set IE=-10 CA=-69

Acquisition tests

6:40 attempting to repeat last night guiding failures, this time with no large field gradient (Moon not yet risen). At the same time will see what the initial WFE is as we preset to stars around the field, should be indicative of ~ 200% of the error we expect to see between off-axis field point and science focal plane. Collimating on axis (wfsc 9-26 ) (note the slightly higher than usual Z12-21 and the large initial Z22, once Z22 corrected total forces reported in PSF GUI ~75, quite high), then going off-axis to GS 16

6:55 acquired GS 16 (acquire 4.fits) collimating wfsc 30-33 preset to GS 30

6:59 acquired GS 30 (acquire 5.fits) collimating wfsc 34-43,

7:07 send preset to going to GS 34 (acquire 6.fits) collimating wfsc 44-46

7:11 sending preset to GS 33 (acquire 7.fits) collimating wfsc 47-52

7:18 sending preset to GS 24 (acquire 8.fits), collimating 53-62 faint and through thin cloud

7:26 sending preset to GS 5 (acquire 9) too many stars, guiding of the faintest one, going to GS 3

7:35 ok, have been around the field, now going to GS 16 in acquire mode and will offset telescope to put GS in corners of field and see if guide preset succeeds.

7:39 IE -> 0, resend preset in Guide mode (acquire 11) acquisition failed "transfer of initial correction to PCS failed", this happened again (acquire 12.fits).

7:41 Wayne "holds" mount and we try again (acquire 13.fits). fails again.

7:43 modify IE -20 and resend preset (acquire14. fits) acquired and started guiding from what was thought to be the "BAD CORNER" last night (acquire 14).

7:49 resent preset as active to collimate, confirm from brightness on WFSC image 71 that we have the brighter star in acquisition image 15. collimating wfsc 71-75

7:53 resend preset in acquire mode and adjust CA to -52 resend preset in Active (acquire 15.fits) acquired GS

7:55 adjust IE=-10 CA=-44 resending preset in ACTIVE mode. (acquire 16.fits) acquisition succedded.

Jesper data for AGw calibrations

8:05 going to take some data for Jesper. sending preset to ::: testing Michelle's change (letting PCS wait longer to make its move for 1830) seems to work, acquired star in 31 seconds, previously would have timed out and failed in 21 sec.

8:18 presetting to BS9107 in Acquire mode to adjust IE CA prior to taking Jesper's data IE=-20 CA=-69

8:21 going to Jesper's astrometric field (IC166) to take data with different position angles, collimating wfsc 85- 88, GS a little faint, resending to GS 1 preset failed couldn't find star though bright enough star seemed to be there. Try GS 2 failed couldnt find star acquire 21.fits Try GS3 succeed collimate wfsc 90-

8:29 Look at IRTC image, see stars with no filter at 1s exposure, trying to find a better guide star to guide off, GS 4 ok. PA 299.17

8:54 taking IRTC data Jesper000110_cube.fit exp 1 s cube of 30 images

8:57 resending preset in TRACK with PA = 299.17 + 45 = 344.17 taking new cube Jesper000111_cube.fit changed exp to 0.2 s due to bright star turning up

9:08 collimating through cloud, not so good, can't take last set of data with collimation resending preset in Track mode taking sequence Jesper000112_cube.fit with PA 254.87 (note PA should have been 254.17 to be 45 degrees off the original PA, but was 254.87 instead).

9:15 now sending preset on-axis in acquire mode (to GS0) with PA 299.17 taking images Guider00001-4 in AGW_Data/081115g

9:21 sending preset to PA 344.17, Guider images 5-8.

9:34 resending preset to PA 254.87. Note that sending presets in Guide mode then killing guiding so there may be a guide correction jump. Guider images 9-12.

9:36 going to take background for IRTC images, need two backgrounds, one fo rJesper 1 with exp 1s 30 frames, next for Jesper 2,3 with 0.2 s, 150 frames. Offset 120 0

9:40 took first Jesper BG file but had big ugly star in the middle of it, so offset 240 " (was 120") and tried again, so JesperBG00114_cube.fit is the 1 s 30 frame one. Jesper BG 00115_cube.fit is the 0.2 s 150 frame one

9:45 preset to BS9107 confirm impression that seeing is >5" (Polaris says 7") turning over to Nancy to play with Lucifer offsets.

LUCIFER preset and offset tests

10:45 John updated the Lucifer pointing model to be the IRTC one but we forgot to restart PCS to force the change (authorizing the instrument is not sufficient), so after doing the IE CA reset we went back restarted PCS and found the star on the hotspot at the original IE CA. Nancy now sending preset, star horrible (seeing >5").

11:01 look at wfsc 130-146 as an example of collimation in spectacularly bad seeinp. Initial numbers were bad and now seems to have converged to < 400 nm !!!

The LUCIFER observing scripts were partially disable by Volker, so we could only send simple presets and offsets from the LUCIFER interface. Therefore we switch to IRAF to try to exercise the dither cadences.

11:05 testing Dave's new dithering algorithm "dither_seq.cl", but had to restart IRAF and load the new LBTtools package in order to find it.

11:12 cl < IRTC$dither_seq.cl
> Problem with SetIRParam command. res=1

11:15 cl < IRTC$dither_seq.cl
> Problem with SetIRParam command. res=1

11:18 Reboot IRTC Windows machine. IRTC software didn't start correctly the first time. I had to kill and restart it.

11:31 cl < IRTC$dither_seq.cl
Something in this sequence caused us to lose the guidestar. (but we never tracked down what the problem was)

11:35 something in the IRAF window is highly confused as it doesn't know the pwd command in the pointto script.

11:41 pointto M5_0881 gs=0 ACTIVE, but this one is too low (42 deg) and too bright.

11:43 pointto BS9108 gs=0 ACTIVE

11:51 cl < IRTC$dither_seq.cl (a correct failure, but not very good reporting up the chain)

Fri Nov 14 23:51:19.022 2008  IIF IIF Offset command started for left LUCIFER from IIF
Fri Nov 14 23:51:22.386 2008  IIF IIF Offset failure for left LUCIFER: oacontrol indicated an error moving to new position
Fri Nov 14 23:51:22.505 2008  IIF IIF Offset command failed for left LUCIFER from IIF

Nov 14 23:51:22 lbtmu05 irc[15470]: OffsetPoiting ANGLE=0. OffRA=5. OffDEC=7. Sys=RADEC OPE=MOUNT NP=true Move=REL Side=left
Nov 14 23:51:22 lbtmu05 irc[15470]: OffsetPointing result status: Error
Nov 14 23:51:22 lbtmu05 irc[15470]: oacontrol indicated an error moving to new position

Nov 14 23:51:22 lbtmu107 LBT_left_GCS: [E] AIP AGW unit error: AGW1: Requested position is Out of Range
Nov 14 23:51:22 lbtmu107 LBT_left_GCS: AGW::setxy(-3.0, 616.2) returned 255
Nov 14 23:51:22 lbtmu107 LBT_left_GCS: error: setProbePosition: couldn't move probe to position (SFP) (x,y,z): 3.028327, -4.003020, 31.570542 mm.
Nov 14 23:51:22 lbtmu107 LBT_left_GCS: offsetGuiding command failed as probe couldn't be moved on new position.

11:58 pointto BS9108 gs=21 ACTIVE and take IRTC out of video mode.

00:03 cl < IRTC$dither_seq.cl

This dithering is working now, but IRTC is having some sort of directory error writing sequences to /Repository/IRTC_Data/081115/. This is because I put in Linux slashes and not windows slashes.

00:14 cl < IRTC$dither_seq.cl

Still got a directory creation error with the opposite slashes. ???? I think the remote mounted directory must be confused. I don't know if the cubes are being saved anywhere or not. We get this error even if the Path FITS is r:.

Repository on 'NFS server (Samba 3.0.28-0.fc7)(nfs.lbt.as.arizona.edu)' (R:)

Field aberration data

1:01 we are unable to reconnect IRTC to the network, so for now will save IRTC cubes to C:/IRTC_Data 081115 on the IRTC computer itself, hopefully this problem will be resolved tomorrow. Seeing has improved to 2", so by a factor of 2. Cleared all active optics and secondary offloads and C00, now using primary mirror for active optics

1:15 seeing varying between 1.5 and many arcseconds. Reset IE=-10 CA=-64 .

1:20 send to BS9107 going to take some field data points, collimateing on axis, measuring off-axis, prior to doing EF.

1:21 collimating in idl on GS 0 wfsc 2- 8,

1:30 GS 35 taking data wfsc 10-12 DA = -212, Y = 464 Back on AXIS (BOA)

1:31 collimating on axis wfsc 13-16

1:41 GS 82 wfsc 17-19 DA = -287 Y = 485.6 (BOA)

1:44 collimating on axis wfsc 20-25

1:49 GS 111 wfsc 26-28 DA -337, Y = 413.6 (BOA)

1:52 collimating wfsc 29-30

1:54 GS 92 failed to acquire, BOA

1:56 collimating wfsc 31-36

1:57 GS 97 wfsc 38-40 DA= -387 Y = 440 (seeing degrading to > 2") (BOA)

2:02 collimating wfsc 41-42

2:04 GS 64 wfsc 43-45 DA = -443 Y =446 (BOA)

2:08 collimating wfsc 46-47

2:09 GS 19 wfsc 48-50, DA = -492 Y = 477 (BOA)

2:12 collimating wfsc 51-53

2:15 GS 8 wfsc 54-56 DA =-181 Y = 432

Now, collimating off axis and taking IRTC images of the in-focus on-axis star, checking for visible aberration in the in-focus image, variations in focus etc. but seeing pretty variable...

2:40 reset IE CA to -10 -64

2:45 taking data collimating on GS 97 and infocus data on IRTC, saved to C:/IRTC_Data 081115 on the IRTC computer (97_infocus0003-5) (ignore frame 3)

Norm and Nancy managed to get the Repository connection back up, so data now moved back to /Repository/IRTC_Data/081115. See Issue #1839.

3:10 collimating on GS 64, GCS wfsc 308--and taking IRTC data 64_infocus0006-7

3:26 collimating on GS 11, collimating to gcs wfsc 325 taking IRTC data 11_infocus0009-10

3:35 collimating on GS 35, collimating gcs wfs 326-327 taking IRTC data 35_infocus0011-12

More Jesper data on M67

3:45 going to M67 GS 4 to collimate and take data for Jesper

3:55 nice and centered on M67 A-D and seeing suddenly plummets, waiting it out...

4:00 taking data cube M67_287.26_00013.fits (PA = 287.26)

4:03 sending preset to GS 4 with PA 247.26 preset fails to find guidestar

4:05 going back to PA indef, collimating then will send Track preset to target with new PA.

4:08 taking cube M67_247.26_00014.fits

4:09 going back to GS 4 PA indef, collimating then will send Track preset to target with new PA 207.26

4:12 taking cube M67_207.26_00015.fits

4:15 going back to GS 4 PA indef, collimating then will send Track preset to target with new PA 167.26

4:19 taking cube M67_167.26_00016.fits

4:21 going back to GS 4 PA indef, collimating then will send Track preset to target with new PA 127.26

4:25 taking cube M67_127.26_00017.fits

4:26 going back to GS 4 PA indef, collimating then will send Track preset to target with new PA 87.26

4:29 taking cube M67_87.26_00018.fits

4:34 going back to GS 4 PA indef, collimating then will send Track preset to target with new PA 47.26

4:37 taking cube M67_47.26_00019.fits

4:37 no backgrounds taken because these are all rotated and can subtract from each other

4:39 going to M5_0258 for EF data, collimating off-axis, EF on axis...

4:50 all set up collimating on GS 4 taking EF00020-21.fits first time ran home$inex_seq.cl and it ran through the motions of defocussing the images but had left the camera on video mode so screwed up, start again...

5:00 collimating on GS 4 wfsc 356-357 taking EF data EF00020-21.fits right at the end OSS crashed and M2 failed to refocus but I think the data is alright

5:06 went to GS 3 collimating 358-359 taking EF data EF00022-23.fits

5:13 collimating on GS 10 wfsc 360-361 taking EF data EF00024-25.fits

5:20 collimating on GS 9 wfsc 362-363 taking EF data EF00026-27.fits

5:27 collimating on GS 7 wfsc 364-366 taking EF data EF00028-29.fits

5:33 taking BG for EF. EFBG00030.fits

5:37 Going to BS 9122 to check WFE changes with offsets as claimed by Nancy. Plan is to collimate do an offset measure wfe do an offset measure wfe etc. then repeat with same cadence without doing the offsets. sending active preset to GS 45

5:39 collimating in IDL on GS 45 wfsc failed to acquire star too faint but catalog shows it nie and bright, trying for GS 37...

5:47 collimating on GS 47 in IDL wfsc 58 sending 10" RA offset measuring wfsc 60-62 sending -10" RA offset measuring wfsc 63-65 offset Dec 10" wfsc 66-68 offset Dec -10" wfsc 69-71 so we see no degradation in WFE with offset in the shack hartman, repeating the experiment with IRTC images of star after collimating, then doing offsets as above (is vibration making a percieved decrease in IQ that SH is insensitive to?

6:15 taken offset 10 0, -10 0, 0 10, 0 -10, but IQ probably not good as sun too bright to collimate should repeat this experiment tomorrow night with more time to get better data.

signing off. IE=-18 CA=-64

-- AndrewRakich - 15 Nov 2008
Topic revision: r6 - 16 Nov 2008, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback